# Efficient Artifacts Removal for Adaptive Deep Brain Stimulation and a Temporal Event Localization Analysis

### v1

@@ 5. Discussion
Stimulation-induced artifacts, including stimulus and transient DC components, pose major challenges for aDBS implementation. Effective artifact removal is critical for accurate neural signal detection and reliable feedback control. We proposed SMARTA+, a back-end enhancement of SMARTA [24], aimed at improving computational efficiency and transient DC artifact suppression while maintaining artifact removal performance. Key improvements include replacing the KNN search with ANN and incorporating wavelet-based dimension reduction, reducing computation time to under 10% of that required by SMARTA. SMARTA+ achieved comparable artifact suppression (as measured by AR and SC) and demonstrated improved recovery of clean signals, reflected in lower NMSE over different spectral bands. temporal event localization analysis further confirmed its effectiveness in accurately detecting beta burst events. To our knowledge, this is the first study to evaluate how accurately a stimulus artifact removal algorithm can enhance beta burst detection.
@@
We now discuss how SMARTA+ achieves the encouraging NMSE results. Although the $\mathrm{NMSE}_{\beta}$ for the semi-real aDBS signals was as low as -2.8 shown in Table 3, reflecting the benefits of carefully chosen sampling rates and stimulation frequencies, the high recall but low precision in beta burst detection from raw signals indicates limited effectiveness. Stimulation artifacts might still mislead aDBS algorithms, causing frequent false detections despite the optimized signal design for PD therapy. This issue arises mainly from two factors. First, while the sampling rate and stimulation frequency have been carefully designed to avoid aliasing, the time-varying morphology of stimulus artifacts still introduces spectral leakage via aliasing. To appreciate this statement, consider the following simple model. Suppose the stimulus artifacts are modeled as $f(t)=A(t)\sum_{j\in\mathbb{Z}}s(t-t_{j})$, where $s$ is smooth and compactly supported modeling the morphology of stimulus artifact so that $s(t)=\alpha_{0}+\sum_{k=1}^{\infty}\alpha_{k}\cos(2\pi kt+\beta_{k})$, $\alpha_{0}\in\mathbb{R}$ is the mean value of $s$, $\alpha_{k}\geq 0$ and $\beta_{k}\in[0,2\pi)$, $t_{j}=j/130$ is the timing of stimulation and $A(t)$ is a positive smooth function modeling the strength of stimulus artifact. As discussed in [24], due to the soft matter property of brain, in general the stimulus artifact changes from time to time, particularly the strength. Despite its simple behavior, we know that $\hat{f}(\xi)=\alpha_{0}\hat{A}(\xi)+\frac{1}{2}\sum_{k=1}^{\infty}\alpha_{k}[e^{\beta_{k}}\hat{A}(\xi-130k)+e^{-\beta_{k}}\hat{A}(\xi+130k)]$, where the equality holds in the uniformly. Therefore, when the magnitude $A$ is large, unless $A$ is a constant function or the support of $\hat{A}$ is compact and “very” concentrated around 0, the spectral spreading of $\hat{A}$ could contaminate the beta band via aliasing. Second, abrupt stimulus onsets and transient DC artifacts introduce broadband distortions that overlap with true beta activity, inflating beta amplitude and triggering false detections. This can be specifically seen in the improvement of NMSEα shown in Table 3. Artifact removal using template subtraction, pulse blanking, or SMARTA improves precision and F1-score by reducing these false positives. However, residual transient artifacts, especially at stimulation onset, remain challenging. SMARTA+ addresses this by also targeting onset-related DC transients, resulting in further gains in precision and F1-score, especially during stimulation onset.
@@
A directly related topic is the existing aDBS implementations. On the Summit RC+S system (Medtronic), aDBS is challenged by both stimulus artifacts and DC transients [1]. While front-end blanking effectively removes stimulus artifacts, residual DC transients introduce broadband noise that can degrade beta burst detection and lead to erroneous self-triggering. In [1], this was mitigated by back-end detector blanking with a 550 ms window. However, such a long blanking duration risks omitting relevant beta activity shortly after stimulation, potentially impairing aDBS performance. An alternative approach was explored in [37], where LFPs recorded using the Percept PC system were analyzed. By carefully aligning the stimulation frequency with the sampling rate, stimulus artifacts were minimized within the beta band. However, this study did not implement aDBS or address DC transient artifacts, leaving it unclear whether self-triggering would still occur under aDBS deployment. Also, the spectral leakage via aliasing due to time-varying stimulus artifact morphology is also inevitable in this situation. As a back-end method, SMARTA+ is adaptable to a range of systems, including Summit RC+S and Percept PC, for removing both stimulus and transient artifacts. Given the repeated stimulation epochs in aDBS signals, SMARTA+’s ability to suppress artifacts at each onset supports more reliable beta detection and offers promise for improving adaptive stimulation control.
@@
The amplitude of transient artifacts has been shown to correlate with the rate of change in stimulation amplitude [12]. To mitigate these artifacts, one strategy involves modulating the stimulation amplitude between a predefined upper and lower limit, rather than turning the stimulation fully on or off [40]. This controlled modulation reduces the rate of change and, consequently, diminishes the magnitude of transient artifacts. Nevertheless, despite the reduction in artifact severity, both transient DC artifacts and stimulus artifacts are not entirely eliminated. Furthermore, the effectiveness of this approach is highly dependent on the selection of the lower amplitude limit. In certain cases, this lower bound may approach zero, at which point the resulting transient artifacts can still be substantial. Therefore, even though dual-threshold aDBS algorithms contribute to artifact reduction, the use of SMARTA+ remains beneficial as an additional step to further improve signal quality by suppressing remaining artifacts.
In aDBS applications, stimulation parameters are often restricted to reduce artifact contamination. A typical example is the above-mentioned frequency planning [37] to prevent aliasing peaks from overlapping with the beta band. While cDBS allows more flexibility in tuning stimulation frequency, such options are further limited in aDBS unless artifacts are effectively removed. SMARTA+ lifts this constraint by reliably eliminating stimulus artifacts, thereby allowing broader frequency choices. Another constraint lies in electrode selection. To achieve common-mode rejection, stimulation contact is typically positioned between two recording contacts, limiting viable configurations in clinical practice. Although directional electrodes are increasingly used in DBS to achieve targeted stimulation [38], their use in aDBS has been limited by the necessity of common-mode rejection. Although SMARTA+ was applied to bipolar recordings in this study, it is also effective for unipolar signals. This versatility enables the possibility of integration of directional electrodes into aDBS systems while maintaining robust artifact suppression.
@@
Beta-band power in LFPs is commonly used as the feedback signal for aDBS in Parkinson’s disease. However, relying solely on beta activity may not fully capture the neural dynamics underlying motor and non-motor symptoms [16]. Gamma-band activity, for instance, correlates with movement and motor performance [36, 25], while the alpha band has been linked to cognitive and emotional processing [27], highlighting the potential value of multi-band biomarkers. Yet, incorporating multiple frequency features into aDBS is challenging when stimulation artifacts are not properly addressed. Without effective artifact removal, stimulation frequencies must be carefully chosen to avoid overlap with aliasing peaks across all bands of interest, restricting biomarker flexibility. By suppressing artifact contamination across the full frequency spectrum, SMARTA+ enables more accurate LFP recovery and supports the integration of diverse biomarkers and paves the way for more precise and adaptive aDBS strategies. This will be explored in our future research.
@@
Adaptive DBS has been investigated for conditions beyond PD, using control biomarkers derived from signals other than subthalamic LFPs. In essential tremor, for example, low-frequency oscillations recorded from subdural cortical electrodes over the primary motor cortex and from the ventral intermediate nucleus have been used for aDBS control [29]. While clinical trials for cognitive or psychiatric applications of aDBS are still lacking, neurophysiological studies suggest strong potential in these areas [10]. Such applications may require diverse biomarkers from multiple brain regions and potentially different recording modalities. A key advantage of SMARTA+ is its versatility. It is not restricted to LFPs and can be adapted to various signal types, given appropriate training data. This flexibility positions SMARTA+ as a potential tool for artifact removal across a range of aDBS research settings, enabling the development of more generalizable and disorder-specific aDBS systems. For example, VHFOs (>> 1000 Hz) in EEG signals are relevant to epilepsy [39] and can serve as potential biomarkers. However, these VHFOs can be contaminated by stimulus artifacts if not properly processed. SMARTA+ has a potential to be applied to remove these artifacts, aiding in the development of aDBS systems for the treatment of epilepsy.
@@
This work has several limitations. While SMARTA+ is substantially faster than SMARTA, it remains computationally intensive for real-time use compared with traditional back-end template subtraction or blanking algorithms, particularly when considering battery constraints. Addressing this will require not only advances in energy-efficient hardware but also further algorithmic optimization. Future directions include incorporating randomized algorithms [17, 18] and exploring cloud-based processing. As an algorithm-focused study, we did not explore leveraging hardware properties to further optimize SMARTA+. Such considerations are essential for clinical implementation but lie beyond the current scope and will be addressed in future work. The study population is relatively small and limited to Asian subjects from a single hospital. Larger-scale studies are needed to further validate the proposed algorithm.
@@ 6. Conclusion
The SMARTA+ algorithm is proposed as an enhanced and time-efficient extension of SMARTA, with demonstrated effectiveness in removing artifacts in LFP signals recorded during electrical stimulation. For each artifact, an accurate template is constructed by identifying similar segments from both the target signal and a pre-recorded dataset using an approximate nearest neighbors algorithm, guided by a similarity metric based on optimal shrinkage for noise reduction. The underlying LFP signal is then recovered by subtracting this template. Additionally, DC transient artifacts are estimated through a two-stage process involving signal smoothing and polynomial fitting and are subtracted from the signal. Temporal event localization analysis shows that SMARTA+ improves the precision of beta-event detection and reduces the timing deviation of detected events. With a computational cost comparable to the duration of each artifact segment, SMARTA+ represents a promising approach for artifact removal in real-time aDBS algorithms.
@@ 7. Ethics statement
This study involving human participants were reviewed and approved by Chang Gung Medical Foundation Institutional Review Board (CGMH-IRB No.202500934B0). The patients/participants provided their written informed consent to participate in this study. Written informed consent was obtained from the individual(s) for the publication of any potentially identifiable images or data included in this article. The research was conducted in accordance with the principles embodied in the Declaration of Helsinki.
@@ 8. Funding
This work was supported by the National Science and Technology Council, Taiwan (NSTC112-2314-B-182-042-MY3) and the National Health Research Institutes, Taiwan (NHRI-EX113-11104NI).
@@
After segmentation, the LFPs in the matrix $\operatorname{\boldsymbol{X}}$ were removed using the extended optimal shrinkage (eOptShrink) algorithm [34]. For the sake of completeness, we summarize eOptShrink here. Without loss of generality, assume $p\leq n$ and denote $\beta=p/n$. Denote the singular value decomposition (SVD) of $\operatorname{\boldsymbol{X}}$ as
@@
To accelerate the eOptShrink algorithm, the randomization technique described in [11] was applied.

### v2

@@ 5. Discussion
Stimulation-induced artifacts, including stimulus and transient DC components, pose major challenges for aDBS implementation. Effective artifact removal is critical for accurate online neural signal detection and reliable feedback control. We proposed SMARTA+, a back-end enhancement of SMARTA [liu2024], aimed at improving computational efficiency for real-time implementation and transient DC artifact suppression while maintaining artifact-removal performance. Key improvements include replacing the KNN search with ANN and incorporating wavelet-based dimension reduction, reducing computation time to under 10% of that required by SMARTA. SMARTA+ achieved comparable artifact suppression (as measured by AR and SC) and demonstrated improved recovery of clean signals, reflected in lower NMSE over different spectral bands. The temporal event localization analysis further confirmed its effectiveness in accurately detecting beta-burst events. To our knowledge, this is the first study to evaluate how accurately a stimulus artifact removal algorithm can enhance beta burst detection.
@@
We now discuss why SMARTA+ achieves the encouraging NMSE results. Although the $\mathrm{NMSE}_{\beta}$ for the semi-real aDBS signals was as low as -3.4, as shown in Table 3, reflecting the benefits of carefully chosen sampling rates and stimulation frequencies, the high recall but low precision in beta burst detection from raw signals indicates limited effectiveness. Stimulation artifacts might still mislead aDBS algorithms, causing frequent false detections despite the optimized signal design for PD therapy. This issue arises mainly from two factors. First, while the sampling rate and stimulation frequency have been carefully designed to avoid aliasing, the time-varying morphology of stimulus artifacts still introduces spectral leakage via aliasing. To appreciate this statement, consider the following simple model. Suppose the stimulus artifacts are modeled as $f(t)=A(t)\sum_{j\in\mathbb{Z}}s(t-t_{j})$, where $s$ is smooth and compactly supported modeling the morphology of stimulus artifact so that $s(t)=\alpha_{0}+\sum_{k=1}^{\infty}\alpha_{k}\cos(2\pi kt+\beta_{k})$, $\alpha_{0}\in\mathbb{R}$ is the mean value of $s$, $\alpha_{k}\geq 0$ and $\beta_{k}\in[0,2\pi)$, $t_{j}=j/130$ is the timing of stimulation and $A(t)$ is a positive smooth function modeling the strength of stimulus artifact. As discussed in [liu2024], due to the soft-matter properties of the brain, the stimulus artifact generally changes from time to time, particularly the strength. Despite its simple behavior, we know that $\hat{f}(\xi)=\alpha_{0}\hat{A}(\xi)+\frac{1}{2}\sum_{k=1}^{\infty}\alpha_{k}[e^{\beta_{k}}\hat{A}(\xi-130k)+e^{-\beta_{k}}\hat{A}(\xi+130k)]$, where the equality holds uniformly. Therefore, when $A$ is large, unless $A$ is constant or the support of $\hat{A}$ is compact and “very” concentrated around 0, the spectral spreading of $\hat{A}$ could contaminate the beta band. Second, abrupt stimulus onsets and transient DC artifacts introduce broadband distortions that overlap with true beta activity, inflating beta amplitude and triggering false detections. This can be specifically seen in the improvement of NMSEα shown in Table 3. Artifact removal using template subtraction, pulse blanking, or SMARTA improves precision and F1-score by reducing these false positives. However, residual transient artifacts, especially at stimulation onset, remain challenging. SMARTA+ addresses this by also targeting onset-related DC transients, resulting in further gains in precision and F1-score, especially during stimulation onset.
@@
A directly related topic is the existing aDBS implementations. On the Summit RC+S system (Medtronic), aDBS is challenged by both stimulus artifacts and DC transients [anso2022artifactRemoval]. While front-end blanking effectively removes stimulus artifacts, residual DC transients introduce broadband noise that can degrade beta burst detection and lead to erroneous self-triggering. In [anso2022artifactRemoval], this was mitigated by back-end detector blanking with a 550 ms window. However, such a long blanking duration risks omitting relevant beta activity shortly after stimulation, potentially impairing aDBS performance. An alternative approach was explored in [thenaisie2021percept], where LFPs recorded using the Percept PC system were analyzed. By carefully aligning the stimulation frequency with the sampling rate, stimulus artifacts were minimized within the beta band. However, the study [thenaisie2021percept] did not implement aDBS or address DC transient artifacts, leaving it unclear whether self-triggering would still occur under aDBS deployment. Also, the spectral leakage via aliasing due to time-varying stimulus artifact morphology is also inevitable in this situation.
As a back-end method, SMARTA+ is adaptable to various neuromodulation platforms, including Summit RC+S and Percept PC, for real-time removal of both stimulus and transient artifacts. The initial decision trees used by SMARTA+ can be pre-built, with updates during stimulation-off periods when the intensity ramps down and the aDBS controller continues collecting samples for online beta-amplitude estimation [little2013aDBS, tinkhauser2017aDBS]. As shown in Tables 1 and 2, SMARTA+ processes each artifact segment faster than the stimulation interval at a 130-Hz sampling rate. These results indicate that SMARTA+ is well suited for real-time aDBS deployment, where its ability to suppress artifacts at each stimulation onset enables more reliable beta detection and may enhance adaptive stimulation control.
@@
The amplitude of transient artifacts is known to correlate with the rate of change in stimulation intensity [hammer2022deltaStimulation]. A common mitigation strategy is to modulate the stimulation between predefined upper and lower bounds instead of switching it fully on or off [velisar2019aDBS]. This controlled ramping reduces the rate of change and thereby attenuates transient artifacts. Nevertheless, despite the reduction in artifact severity, both transient DC artifacts and stimulus artifacts are not entirely eliminated, and the degree of reduction depends critically on the chosen lower amplitude limit. In certain cases, this lower bound may approach zero, yielding substantial artifacts despite dual-threshold control. Thus, while dual-threshold aDBS helps reduce artifact severity, applying SMARTA+ as an additional processing step remains valuable for further improving signal quality.
Artifact contamination also constrains stimulation parameters in aDBS. For example, the above-mentioned frequency planning is often required to prevent aliasing peaks from overlapping with the beta band [thenaisie2021percept]. Although cDBS allows broader frequency tuning, aDBS is more restricted unless artifacts are effectively suppressed. By reliably removing stimulus artifacts, SMARTA+ relaxes these limitations and enables more flexible frequency selection.
Electrode configuration presents another challenge. Common-mode rejection typically requires placing the stimulation contact between two recording contacts, restricting viable electrode arrangements in clinical practice. Although directional electrodes offer improved targeting [tinkhauser2018directional], their use in aDBS has been limited by the need for common-mode rejection. While SMARTA+ was demonstrated here using bipolar recordings, it is also effective for unipolar signals, thereby facilitating the integration of directional electrodes into aDBS systems while maintaining robust artifact suppression.
@@
Beta-band LFP power is the most widely used control biomarker for aDBS in Parkinson’s disease, yet beta activity alone may not fully represent the underlying neural state [johnson2016beyondbeta]. Gamma-band activity correlates with movement and motor performance [swann2018aDBSgamma, lofredi2018aDBSgamma], and alpha band dynamics are linked to cognitive and emotional processing [marceglia2011alpha], motivating the use of multi-band biomarkers. However, incorporating multiple frequency features becomes difficult when stimulation artifacts are not adequately addressed, as stimulation frequencies must then be chosen to avoid aliasing across all bands of interest. By suppressing artifacts across the entire spectrum, SMARTA+ enables more accurate recovery of LFP activity, supports the use of diverse biomarkers, and paves the way for more flexible and precise aDBS strategies, an avenue we will explore in future work.
@@
Adaptive DBS is also being investigated for disorders beyond PD, using control biomarkers derived from signals other than subthalamic LFPs. In essential tremor, for example, low-frequency oscillations recorded from subdural cortical electrodes over the primary motor cortex and from the ventral intermediate nucleus have been used for aDBS control [opri2020ET]. While clinical aDBS trials for cognitive or psychiatric applications remain limited, neurophysiological studies indicate strong potential [guidetti2021aDBSintro]. Such applications may require diverse biomarkers across multiple brain regions and recording modalities. A key advantage of SMARTA+ is its versatility. Theoretically, it is not restricted to LFPs and can be adapted to different signal types, sampling rates, and stimulation frequencies, as partially demonstrated in Sections 4.5 and 4.6. This flexibility positions SMARTA+ as a promising artifact-removal tool across a broad range of aDBS research settings, supporting the development of generalizable and disorder-specific adaptive stimulation systems. For example, VHFOs (>> 1000 Hz) in EEG signals are relevant to epilepsy [usui2015VHFO] and can serve as potential biomarkers. However, these VHFOs can be contaminated by stimulus artifacts if not properly processed. Applying SMARTA+ to such signals may facilitate artifact-free VHFO extraction and contribute to advancing aDBS approaches for epilepsy.
@@
This work has several limitations. The dedicated hardware required to implement the front-end approaches described in [culaclii2016ts, stanslaski2018aDBSdevice, anso2022artifactRemoval] was not available in this study; therefore, these methods were simulated using software-based implementations. As a result, their performance may not fully reflect the optimal capabilities of the original hardware systems. Nevertheless, SMARTA+ was evaluated under identical conditions, without incorporating any front-end artifact removal processes. While SMARTA+ is substantially faster than SMARTA, it remains computationally intensive compared with traditional back-end template subtraction or blanking algorithms, particularly when considering battery constraints. Addressing this will require not only advances in energy-efficient hardware but also further algorithmic optimization. Future directions include incorporating randomized algorithms [jones2011RANN, jones2013RANN] and exploring cloud-based processing. As an algorithm-focused study, we did not explore leveraging hardware properties to further optimize SMARTA+. Such considerations are essential for clinical implementation but lie beyond the current scope and will be addressed in future work. The study population is relatively small and limited to Asian subjects from a single hospital. Larger-scale studies are needed to further validate the proposed algorithm.
@@ 6. Conclusion
The SMARTA+ algorithm is proposed as an enhanced and time-efficient extension of SMARTA, with demonstrated effectiveness in removing artifacts in LFP signals recorded during electrical stimulation. For each artifact, an accurate template is constructed by identifying similar segments from both the target signal and a pre-recorded dataset using an approximate nearest neighbors algorithm, guided by a similarity metric based on optimal shrinkage for noise reduction. The underlying LFP signal is then recovered by subtracting this template. Additionally, DC transient artifacts are estimated through a two-stage process involving signal smoothing and polynomial fitting and are subtracted from the signal. Temporal event localization analysis shows that SMARTA+ improves the precision of beta-event detection and reduces the timing deviation of detected events. With a computational cost lower than the duration of each artifact segment, SMARTA+ demonstrates suitability for integration into real-time aDBS algorithms for artifact removal.
@@ 7. Ethics statement
This study was reviewed and approved by the Chang Gung Medical Foundation Institutional Review Board (CGMH-IRB No. 202500934B0) for secondary analysis of de-identified data obtained from prior IRB-approved studies. Informed consent had been obtained from all participants under the original IRB protocols. The research was conducted in accordance with the principles of the Declaration of Helsinki.
@@ 8. Funding
This work was supported by the National Science and Technology Council, Taiwan (NSTC112-2314-B-182-042-MY3, NSTC112-2115-M-002-015-MY3, NSTC114-2811-M-002-158) and the National Health Research Institutes, Taiwan (NHRI-EX113-11104NI).
@@
After segmentation, the LFPs in the matrix $\operatorname{\boldsymbol{X}}$ were removed using the extended optimal shrinkage (eOptShrink) algorithm [su2022cOS]. For the sake of completeness, we summarize eOptShrink here. Without loss of generality, assume $p\leq n$ and denote $\beta=p/n$. Denote the singular value decomposition (SVD) of $\operatorname{\boldsymbol{X}}$ as
@@
To accelerate the eOptShrink algorithm, the randomization technique described in [halko2011finding] was applied.